# Accelerating Natural Gradient Descent for PINNs with Randomized Numerical Linear Algebra

### v2

@@ 5.2 Heat equation in 3+1D
$\begin{cases}\partial_{t}u-\Delta u=f&I\times\Omega=(0,1)\times(0,1)^{3},\\ u=g_{D}&I\times\partial\Omega,\\ u(0,\cdot)=u_{0}&\Omega,\end{cases}$ (13)
with exact solution $u(t,x,y,z)=\left(\cos(\pi x)+\cos(\pi y)+\cos(\pi z)\right)e^{-\pi^{2}t/4}$ as in [72]. The corresponding least squares loss and metrics are
$E(u)=\frac{1}{2}\left\|{\partial_{t}u-\Delta u-f}\right\|_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\partial\Omega))}^{2}+\frac{1}{2}\left\|{u(0,\cdot)-u_{0}}\right\|_{\mathrm{L}^{2}(\Omega)}^{2},$
${g}\left({u},\,{v}\right)=\left({\partial_{t}u-\Delta u},\,{\partial_{t}v-\Delta v}\right)_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))}+\left({u(0,\cdot)},\,{v(0,\cdot)}\right)_{\mathrm{L}^{2}(\Omega)}.$
Further details are provided in Section E.2.
Also in this example, NyströmNGD provides the best accuracy at a reduced computational cost, as shown in Figure 3. All other methods saturate at a lower accuracy, despite being given a higher computational budget. The Nyström rank parameter $\ell$ rapidly increases and then stabilizes at a value lower than $\ell_{\max}=500$, although at a higher value compared to the Poisson example.
Figure 3: Heat equation in 3+1D discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
@@ 5.2.1 Kovasznay Flow (2D steady state Navier–Stokes flow)
We consider the two-dimensional steady Navier–Stokes flow as originally described in [45], with Reynolds number $\mathrm{Re}=40$, following the setup in [38, Section 4.1]. The system of PDEs reads
$\begin{cases}-\nu\Delta u+(u\cdot\nabla)u+\nabla p=f&\Omega=(-\frac{1}{2},1)\times(-\frac{1}{2},\frac{3}{2}),\\ \nabla\cdot u=0&\Omega,\\ u=g_{D}&\partial\Omega.\end{cases}$ (14)
The nonlinear least squares loss is
$E(u,p)=\frac{1}{2}\left\|{-\nu\Delta u+(u\cdot\nabla)u+\nabla p-f}\right\|^{2}_{\mathrm{L}^{2}(\Omega)^{2}}+\frac{1}{2}\left\|{\nabla\cdot u}\right\|^{2}_{\mathrm{L}^{2}(\Omega)}+\frac{1}{2}\left\|{u-g_{D}}\right\|^{2}_{\mathrm{L}^{2}(\partial\Omega)},$
and as a metric we consider the Gauss–Newton metric (9). We employ two separate neural networks for $u$ and $p$. Further details are provided in Section E.3.
@@
NyströmNGD still provides the best accuracy in fewer than $1000$ iterations, as shown in Figure 4. However, compared to previous examples, the rank $\ell$ grows to the maximum rank $\ell_{\max}=500$ before stabilizing around $400$, indicating that the Gramian is less suited to low-rank approximation than in previous examples. This explains the smaller difference, both in final accuracy and computation time, between NyströmNGD and BFGS. Nevertheless, the gap in iterations remains substantial.
Figure 4: Kovasnay Flow discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations (first $1000$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
5.3 Source recovery problem for the Poisson equation in 3D
We consider the inverse problem of recovering both the solution $u$ and the source term $f$ in the 3D Poisson equation (12) from $N_{\mathrm{data}}=40$ noiseless observations $\{u^{\star}_{i}\}_{i=1}^{N_{\mathrm{data}}}$ of the true solution at random points $\{\mathbf{x}_{i}\}_{i=1}^{N_{\mathrm{data}}}$ in $\Omega$. The corresponding least squares loss and metrics are
$E(u,f)=\frac{1}{2}\left\|{\Delta u+f}\right\|_{\mathrm{L}^{2}(\Omega)}^{2}+\frac{1}{2}\left\|{u-g_{D}}\right\|_{\mathrm{L}^{2}(\partial\Omega)}^{2}+\frac{1}{2N_{\mathrm{data}}}\sum_{i=1}^{N_{\mathrm{data}}}\lvert u(\mathbf{x}_{i})-u^{\star}_{i}\rvert^{2}$
${g}\left({(u,f)},\,{(v,h)}\right)=\left({\Delta u+f},\,{\Delta v+h}\right)_{\mathrm{L}^{2}(\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\partial\Omega)}+\left({u},\,{v}\right)_{\mathrm{L}^{2}(\Omega)}.$
We employ two separate neural networks for $u$ and $f$. Further details are provided in Section E.4.
As shown in Figure 5, NyströmNGD achieves higher accuracy than NGD and NGD with CG, and converges in fewer iterations than BFGS. However, BFGS can perform more iterations within the same time budget and eventually surpasses NyströmNGD in accuracy. The reduced performance of NyströmNGD can be explained by the poor low-rank approximability of the Gramian, as indicated by the rank parameter $\ell$ reaching the threshold $\ell_{\max}=500$. This suggests that $\ell_{\max}$ matrix-vector products are insufficient for accurately approximating the Gramian via randomized Nyström, leading to a less effective preconditioner.
Figure 5: Source recovery problem for the Poisson equation in 3D discretized via PINNs. From left to right: relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations (first $500$ iterations), relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. wall-clock time, and Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
@@ 6 Discussion and Conclusion
We introduced NyströmNGD, a novel optimization algorithm that uses randomized Nyström preconditioning to accelerate the inner conjugate gradient (CG) solver in matrix-free natural gradient descent (NGD). This approach significantly reduces the cost of the inner solver, resulting in faster convergence in both iterations and wall-clock time. Moreover, our work extends matrix-free NGD to a broader class of metrics relevant to PDE-related problems. We validated our method on different PDE problems discretized using PINNs, demonstrating that NyströmNGD consistently outperforms prior NGD methods by achieving higher accuracy at a lower computational cost.
Limitations and future work
The use of randomized Nyström for preconditioning builds on the empirically observed low-rank structure of the Gramian. Our results confirm that in many cases the Gramian admits strong low-rank compression, though this is not always guaranteed and warrants further analysis. Nevertheless, the core idea of preconditioning the inner CG solver remains applicable, and we plan to explore alternative Gramian approximations for use as preconditioners. Notably, NyströmNGD not only reduces the computational cost but also improves final accuracy over standard NGD. This may be due to improved numerical stability or the stochastic nature of the Nyström approximation, which could help escape poor local minima, similar to how SGD can outperform full-batch GD in traditional machine learning [44, 39, 34]. This aspect merits further investigation. Finally, NyströmNGD is especially promising in the context of neural network-based solvers for variational formulations of PDEs, where the choice of metric can be guided by the underlying variational structure. Our preliminary results (see Appendix F) support the potential of the proposed approach and motivate future research in this direction.
Acknowledgments and Disclosure of Funding
@@
Appendix A Pseudocode of Randomized Nyström Approximation
3:$\boldsymbol{\Omega}=\textrm{randn}(p,\ell)$triangleright Gaussian test matrix
[…]
5:$\mathbf{Y}=\mathbf{G}\boldsymbol{\Omega}$triangleright $\ell$ matvecs with $\mathbf{G}$
[…]
7:$\mathbf{Y}_{\nu}=\mathbf{Y}+\nu\boldsymbol{\Omega}$triangleright Shift for stability
8:$\mathbf{C}=\textrm{chol}(\boldsymbol{\Omega}^{\top}\mathbf{Y}_{\nu})$
9:$\mathbf{B}=Y_{\nu}/\mathbf{C}$triangleright Triangular solve
10:$[\mathbf{U},\boldsymbol{\Sigma},\sim]=\textrm{svd}(\mathbf{B},0)$triangleright Thin SVD
11:$\hat{\boldsymbol{\Lambda}}=\max\{0,\boldsymbol{\Sigma}^{2}-\nu\mathbf{I}\}$triangleright Remove shift
Algorithm 2 Randomized Nyström Approximation [26, 51]
Appendix B An extended discussion on the choice of the metric
In this Appendix section, we further elaborate on the selection of the metric discussed in Section 3.1, providing additional details and extending the discussion.
It can be shown that NGD corresponds to Riemannian gradient descent [14] on the manifold $(\mathcal{M},g)$, up to the $g$-orthogonal projection of the Riemannian gradient onto the model’s generalized tangent space $\mathrm{T}_{\boldsymbol{\theta}}\mathcal{N}$, and a second-order correction; see [55, Theorem 1]. More precisely, the NGD update $\boldsymbol{\theta}_{k+1}=\boldsymbol{\theta}_{k}-\alpha_{k}\mathbf{G}(\boldsymbol{\theta}_{k})^{\dagger}\nabla L(\boldsymbol{\theta}_{k})$ satisfies

### v3

@@ 5.1.3 Heat equation in 3+1D
$\begin{cases}\partial_{t}u-\Delta u=f&\text{in }I\times\Omega=(0,1)\times(0,1)^{3},\\ u=g_{D}&\text{on }I\times\partial\Omega,\\ u(0,\cdot)=u_{0}&\text{in }\Omega,\end{cases}$ (16)
with exact solution $u(t,x,y,z)=\big(\cos(\pi x)+\cos(\pi y)+\cos(\pi z)\big)e^{-\pi^{2}t/4}$, as in [85]. We consider Equation˜16 in strong form, with function spaces $\mathrm{V}=\mathrm{L}^{2}(I,\mathrm{H}^{2}(\Omega))\cap\mathrm{H}^{1}(I,\mathrm{L}^{2}(\Omega))$, $\mathrm{X}=\mathrm{L}^{2}(I,\mathrm{L}^{2}(\Omega))$, $\mathrm{Y}=\mathrm{L}^{2}(I,\mathrm{L}^{2}(\partial\Omega))\times\mathrm{L}^{2}(\Omega)$, and operators
$\mathcal{F}(u)=\partial_{t}u-\Delta u,\qquad\mathcal{B}(u)=\big(\left.{u}\right|_{I\times\partial\Omega},\,u(0,\cdot)\big).$
The corresponding least-squares loss reads
with induced metric
Considerations on the coercivity of the metric analogous to Remark˜5.1 apply here; see [85, Section 4.5] for details.
Results
As shown in Figure˜4, NyströmNGD again provides the best accuracy at a reduced computational cost. All other methods saturate at a lower accuracy, despite being given a higher computational budget. Comparing with [85], NGD would not reach the final accuracy of NyströmNGD even when given a larger computational budget of $5000$ iterations. The Nyström rank parameter $\ell$ increases rapidly during early iterations and then stabilizes below $\ell_{\max}=500$, but at a higher value than in the Poisson example, consistent with the slightly slower singular value decay observed in Figure˜5.
(a) Relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations).
(b) Relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time.
(c) Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 4: Convergence for the Heat equation in 3+1D discretized via PINNs.
Figure 5: Heat equation in 3+1D discretized via PINNs and trained with NGD using the SVD-based pseudoinverse. Left: decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. Right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
@@ 5.1.4 Kovasznay Flow (2D steady state Navier–Stokes flow)
We consider the two-dimensional steady Navier–Stokes flow as originally described in [54], with Reynolds number $\mathrm{Re}=40$, following the setup in [46, Section 4.1]. The system of PDEs reads
with exact solution
where $\lambda=\frac{1}{2\nu}-\sqrt{\frac{1}{4\nu^{2}}+4\pi^{2}}$, $\nu=\frac{1}{\mathrm{Re}}=\frac{1}{40}$, and $c$ is chosen such that $\int_{\Omega}p=0$. Following [46], we consider the function spaces $\mathrm{V}=\mathrm{H}^{2}(\Omega)\times\{p\in\mathrm{H}^{1}(\Omega):\int_{\Omega}p=0\}$ and $\mathrm{X}=\mathrm{L}^{2}(\Omega)^{2}\times\mathrm{L}^{2}(\Omega)\times\mathrm{L}^{2}(\partial\Omega)$. Problem (19) is written in strong form as a nonlinear least-squares problem with residual
$\mathcal{R}(u,p)=\begin{bmatrix}-\nu\Delta u+(u\cdot\nabla)u+\nabla p-f\\ \nabla\cdot u\\ u-g_{D}\end{bmatrix}$, (20)
and corresponding energy functional
As a metric, we consider the Gauss–Newton metric
${g_{(u,p)}}\left({(v,h)},\,{(w,t)}\right)=\left({\mathrm{D}\mathcal{R}(u,p)[(v,h)]},\,{\mathrm{D}\mathcal{R}(u,p)[(w,t)]}\right).$
Since the residual (20) is nonlinear, the metric is point-dependent. We refer to [46] for a discussion on coercivity. To discretize (19), we represent $(u,p)=(u_{\boldsymbol{\theta}_{1}},p_{\boldsymbol{\theta}_{2}})$ as two neural networks, with parameters combined into a single vector $\boldsymbol{\theta}=[\boldsymbol{\theta}_{1},\boldsymbol{\theta}_{2}]$.
Results
@@
In this example, NyströmNGD again achieves the best accuracy in fewer than $1000$ iterations, as shown in Figure˜6. However, compared to the previous examples, the Nyström rank parameter $\ell$ exhibits a pronounced peak during intermediate iterations, reaching the maximum rank $\ell_{\max}=500$ before stabilizing around $400$. This behavior indicates that the Gramian is less suited to low-rank approximation in this problem, which explains the smaller differences in final accuracy and computational time between NyströmNGD and BFGS. Nevertheless, NyströmNGD still requires substantially fewer iterations to converge.
(a) Relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations (first $1000$ iterations.)
(b) Relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. wall-clock time.
(c) Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 6: Convergence for the Kovasnay Flow discretized via PINNs.
Figure 7: Kovasnay Flow discretized via PINNs and trained with NGD using SVD-based pseudoinverse. Left: decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. Right: the relative $\mathrm{H}^{1}(\Omega)$ error on $u$ vs. iterations.
5.2 PDEs in weak form: FEINNs
We now consider variational formulations of PDEs using FEINNs, as defined in Section˜2.3.1, to approximate the solution.
Computational details
For all experiments with variational formulations, the parameters in Algorithm˜2 are set as $\ell_{0}=10$, $\ell_{\max}=500$, $\gamma=30$, $\mathtt{maxit}=20$, and $\kappa=0.1$. Neural networks are interpolated onto a finite element space of degree $d=4$ over a uniform quadrilateral mesh with mesh size $h=1/30$, yielding approximately $14,500$ degrees of freedom. Integrals in the loss and metric are evaluated by interpolation onto the FE space and Gauss-Jacobi quadrature of order $2d=8$. Errors are computed with a Gauss-Jacobi quadrature of order $4d=16$. We use the Julia packages Lux.jl [67] for neural networks and Optim.jl [62] for optimization. All experiments are conducted in double precision on a single NVIDIA RTX 4090 GPU with 24 GB of memory.
5.2.1 Poisson problem in 2D – Energy minimization (or Deep Ritz like) formulation
We consider the Poisson problem on the 2D unit square
The solution is then $u=u_{0}+\tilde{g}_{D}$. The discretization spaces are $\mathrm{V}=\mathrm{H}^{1}_{0}(\Omega)$, $\mathrm{X}=\mathrm{V}^{\prime}=\mathrm{H}^{-1}(\Omega)$, with linear operator
$\mathcal{F}:\mathrm{H}^{1}_{0}(\Omega)\to\mathrm{H}^{-1}(\Omega),\qquad\left\langle{\mathcal{F}u},\,{v}\right\rangle_{\mathrm{H}^{-1}\times\mathrm{H}^{1}_{0}}=a(u,v).$
Since $a$ is symmetric and coercive on $\mathrm{V}\times\mathrm{V}$, we adopt an energy minimization formulation, akin to the Deep Ritz method:
$\min_{u_{0}\in\mathrm{H}^{1}_{0}(\Omega)}E(u_{0})\coloneqq\int_{\Omega}\frac{1}{2}\left\|{\nabla u_{0}}\right\|^{2}-fu_{0}+\nabla\tilde{g}_{D}\cdot\nabla u_{0}\,\mathrm{d}\mathbf{x},$ (24)
with corresponding metric $g=a$.
Computational details
We run NGD-based methods for $500$ iterations and BFGS for $1.2\times 10^{5}$ iterations. Since the loss can become negative, we set $\mu=10^{-9}$ for NGD and NGD with CG, and $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-9})$ for NyströmNGD.
Results
NyströmNGD and BFGS achieve comparable final accuracy, as shown in Figure˜8, but NyströmNGD requires a fraction of the iterations and is approximately $10\times$ faster in wall-clock time. This efficiency stems from the strong low-rank approximability of the Gramian, evident in Figure˜9, which leads the Nyström rank parameter $\ell$ to stabilize around $80$, far below the parametric dimension $p$. NGD and NGD with CG plateau at lower accuracy and are not competitive in iterations or runtime.
(a) Relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations).
(b) Relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time.
(c) Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 8: Convergence for the Poisson equation in 2D discretized via FEINNs and energy minimization.
Figure 9: Poisson equation in 2D discretized via FEINNs and energy minimization, trained with NGD using an SVD-based pseudoinverse. Left: decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. Rigth: the relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
5.2.2 Diffusion-reaction-transport problem in 2D
We consider the diffusion-reaction-transport problem
The function spaces are $\mathrm{V}=\mathrm{H}^{1}_{0,\Gamma_{D}}(\Omega)$, $\mathrm{X}=\mathrm{V}^{\prime}=\mathrm{H}^{-1}_{\Gamma_{D}}(\Omega)$, and the variational formulation of (25) is defined as $\mathcal{F}u=f$ where
$\begin{split}\left\langle{\mathcal{F}u},\,{v}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}}&=a(u,v)=\int_{\Omega}(\kappa\nabla u)\cdot\nabla v+(\boldsymbol{\beta}\cdot\nabla u)\,v+\sigma\,u\,v\,\,\mathrm{d}\mathbf{x}\\ \left\langle{f},\,{v}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}}&=\int_{\Omega}F\,v\,\,\mathrm{d}\mathbf{x}+\int_{\Gamma_{N}}\psi\,v\,\,\mathrm{d}\mathbf{s}\end{split}$
Since $\mathcal{F}$ is in general not symmetric positive definite, we adopt the least-squares formulation
$\min_{u\in\mathrm{H}^{1}_{0}(\Omega)}E(u)\coloneqq\frac{1}{2}\left\|{\mathcal{F}u-f}\right\|_{\mathrm{H}^{-1}_{\Gamma_{D}}}^{2}=\left\langle{\mathcal{F}u-f},\,{\mathcal{R}_{\mathrm{V}}^{-1}(\mathcal{F}u-f)}\right\rangle_{\mathrm{H}^{-1}_{\Gamma_{D}}\times\mathrm{H}^{1}_{0,\Gamma_{D}}},$ (26)
with associated metric ${g}\left({v},\,{w}\right)\coloneqq\left\langle{\mathcal{F}w},\,{\mathcal{R}_{\mathrm{V}}^{-1}\mathcal{F}v}\right\rangle$. Here, $\mathcal{R}_{\mathrm{V}}^{-1}=(-\Delta)^{-1}:\mathrm{H}^{-1}_{\Gamma_{D}}\to\mathrm{H}^{1}_{0,\Gamma_{D}}$ denotes the inverse Riesz operator, which involves solving a Poisson equation.
When discretizing (26) and the corresponding metric with FEINNs, applying the discrete Riesz operator $\mathcal{R}_{\mathrm{V}_{h}}^{-1}$ corresponds to solving a Poisson problem with FEM. To reduce computational cost, one could instead use a spectrally equivalent approximation, such as a few cycles of a geometric multigrid preconditioner, as done in [8]. However, for simplicity, we solve the Poisson problem directly using FEM, computing a Cholesky factorization of the Laplacian matrix once.
Computational details
We run NGD-based methods for $1000$ iterations and BFGS for $1.2\times 10^{5}$ iterations. For NGD and NGD with CG we use $\mu=\min(10^{-5},L(\boldsymbol{\theta}))$ as suggested in [85]. For NyströmNGDwe set $\mu=\max(p\cdot\hat{\lambda}_{1}\cdot\epsilon_{\mathrm{mach}},10^{-8}\cdot L(\boldsymbol{\theta})^{2})$.
Results
As illustrated in Figure˜10, NyströmNGD reaches convergence in few iterations and achieves higher accuracy compared to other NGD methods. The gap in convergence speed relative to BFGS is substantial in both iterations and wall-clock time. The smaller difference in wall-clock time between NyströmNGD and BFGS, compared to previous variational examples, is due to the higher Nyström rank parameter $\ell$ at which NyströmNGD stabilizes. Although $\ell$ remains much lower than the parametric dimension $p$, it is higher than in earlier examples, reflecting the slower, yet still significant, decay of the singular values of the Gramian shown in Figure˜11.
(a) Relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations (first $1000$ iterations).
(b) Relative $\mathrm{H}^{1}(\Omega)$ error vs. wall-clock time.
(c) Nyström rank parameter $\ell$ vs. iterations, with $\ell_{\max}=500$.
Figure 10: Convergence for the diffusion-reaction-transport equation (25) in 2D discretized via FEINNs.
Figure 11: Diffusion-reaction-transport equation in 2D discretized via FEINNs, trained with NGD using an SVD-based pseudoinverse. Left: decay of the normalized singular values $\{\sigma_{i}/\sigma_{1}\}_{i=1}^{p}$ of the Gramian across iterations, with each color representing a different iteration. Right: relative $\mathrm{H}^{1}(\Omega)$ error vs. iterations.
@@ 6 Discussion and Conclusion
We introduced NyströmNGD, a novel optimization algorithm that uses randomized Nyström preconditioning to accelerate the inner conjugate gradient (CG) solver in matrix-free natural gradient descent (NGD). This approach significantly reduces the cost of the inner solver, resulting in faster convergence in both iterations and wall-clock time. Moreover, our work extends matrix-free NGD to a broader class of metrics relevant to PDE-related problems. We validate NyströmNGD on various PDEs problems, both in strong form discretized with PINNs and in variational form discretized with FEINNs, demonstrating that it consistently outperforms previous NGD methods by achieving higher accuracy at lower computational cost.
The use of randomized Nyström for preconditioning builds on the empirically observed low-rank structure of the Gramian. Our results confirm that in many cases the Gramian admits strong low-rank compression, though this is not always guaranteed and warrants further analysis. Nevertheless, the core idea of preconditioning the inner CG solver remains applicable, and we plan to explore alternative Gramian approximations for use as preconditioners. Notably, NyströmNGD not only reduces the computational cost but also improves final accuracy over standard NGD. This may be due to improved numerical stability or the stochastic nature of the Nyström approximation. This aspect merits further investigation. Finally, NyströmNGD is especially promising in the context of neural network-based solvers for variational formulations of PDEs, where the choice of metric can be guided by the underlying variational structure. Our results in Section˜5.2 support the potential of the proposed approach and motivate future research in this direction.
CRediT authorship contribution statement
Ivan Bioli: Conceptualization, Methodology, Software, Validation, Formal analysis, Investigation, Data Curation, Writing – Original Draft, Writing – Review & Editing, Visualization
Carlo Marcati: Conceptualization, Methodology, Validation, Formal analysis, Writing – Original Draft, Writing – Review & Editing, Supervision
Giancarlo Sangalli: Conceptualization, Methodology, Validation, Formal analysis, Resources, Writing – Original Draft, Writing – Review & Editing, Supervision, Project administration, Funding acquisition
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
No data were used for the research described in the article. The code used to reproduce the findings of this paper is available from the corresponding author upon request.
Acknowledgments
@@
Appendix A Choice of the Metric
The link with operator preconditioning in Section˜3.2.1 makes it evident that the choice of metric has a crucial influence on the convergence speed of NGD. In some cases, the problem’s structure naturally suggests a metric, as in Variational MonteCarlo [68, 42, 65] or Sobolev gradient flows [41, 19]. In general, the link between NGD and Riemannian gradient descent [64] suggests selecting the metric $g$ to reduce the condition number of the Riemannian Hessian [15]. Indeed, it can be shown that NGD corresponds to Riemannian gradient descent [15] on the manifold $(\mathrm{V},g)$, up to the $g$-orthogonal projection of the Riemannian gradient onto the model’s generalized tangent space $\mathrm{T}_{\boldsymbol{\theta}}\mathcal{N}$, and a second-order correction; see [65, Theorem 1]. More precisely, the NGD update $\boldsymbol{\theta}_{k+1}=\boldsymbol{\theta}_{k}-\alpha_{k}\mathbf{G}(\boldsymbol{\theta}_{k})^{\dagger}\nabla L(\boldsymbol{\theta}_{k})$ satisfies